OpenStudy (anonymous):

which is an equation of the line passing through (3,9) with a slope -2/3?

OpenStudy (anonymous):

point slope formula: \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

\[y-9=-\frac{2}{3}(x-3)\]

OpenStudy (anonymous):

o in understand now thanks so much!

OpenStudy (ashwinram):

Equation of line: y = mx + c. m = 2/3 (known). Plug in known point: 9 = (2/3) * 3 + c ==> solve for c.

OpenStudy (ashwinram):

(or m = -2/3. can't tell due to line break in your question)